OpenStudy (anonymous):

Find the slope and the y-intercept. 3-1/5y=7x

jhonyy9 (jhonyy9):

this want to be 1 3 - --- =7x so than yes ??? ---- will be 15y -1 =35xy --- 15y-35xy =1 --- 5y y(15-35x=1 1 y= -------- 5(3-7x)